How many ways can you pick a president, vice-president, and secretary from a group of six boys and five girls if
a) there are no restrictions?
b) there must be at least one boy chosen?
c) there must be only one girl chosen?
1
Expert's answer
2020-03-09T11:58:23-0400

There are 3 positions. There are 6 boys and 5 girls. Total of 11

a) no restrictions

This is a permutation ;

nPr ="\frac {n!} {(n-r)!}"

=​"\frac {11!}{(11-3)!}"

=990 ways

b) at least one boy chosen.

This is where 1, 2 or 3 boys can be selected.

We subtract when no boy(all girls) is selected from when no restriction is put in place.

11P3-5P3

="\\frac {11!}{(11-3)!}-\\frac {5!}{(5-3)!}\u200b"

=990- 60

=930 ways

c) only one girl chosen

This means that 2 boys are chosen

And there are 3 positions.

5P1*6P2*3

="\frac {5!}{(5-1)!}*\frac {6!}{(6-2)!}*3"

=450 ways